HUBBLE SPACE TELESCOPE DAILY REPORT #2236 PERIOD COVERED: 0000Z (UTC) 10/23/98 - 0000Z (UTC) 10/26/98 Daily Status Report as of 299/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Six Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Six Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Eight Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.4 Completed Four Sets of NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Twenty Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.6 Completed Two Sets of STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a non-scripted parallel POMS test. The observations were completed as scheduled, and no problems were reported. 1.7 Completed Four Sets of NIC/2 7226 (Search for Massive Jupiters) The Near Infrared Camera (Camera Two) was used to make a corographic search for massive planets around the nearby, young main sequence star HD220140 and TWA6. Because of the extreme youth of these objects, any low-mass brown dwarf and planetary companions will still be in a higher luminosity phase and thus easily detectable. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed WF/PC-2 6745 (Fundamental Plane, Morphology-Density Relation, and Lensing in the z=0.58 Arc Cluster CL2053) The WF/PC-2 was used to make observations of the galactic cluster Gal-Clus-2053 in order to study its fundamental plane structure, its morphology-density relation, and its lensing. The distribution of {post}starburst galaxies and/or mergers as a function of radius will be of particular interest. The mass distribution will be determined from the distortion of background galaxies by weak lensing. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed Nine Sets of NIC/2/3 8082 (NICMOS pure parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed WF/PC-2 7431 (Snapshot Survey of Microlensed Source Stars) The WF/PC-2 was used to make snapshot survey images of microlensed source star Macho-95-3. Gravitational microlensing surveys have revealed an excess of lensing events over theoretical predictions toward both the Large Magellanic Cloud and the Galactic Bulge. The observations were completed as planned, and no anomalies were reported. 1.11 Completed FGS 7488 (A Search for Main-Belt Binary Asteroids) The Fine Guidance Sensors were used to make astrometric observations of the main belt asteroid Hector to search for a possible companion moon/asteroid. The targets was chosen on the basis of an extensive analysis based mainly on peculiar photometric properties suggesting a possible binary nature. The observations were completed as planned, and no problems were noted. 1.12 Completed WF/PC-2 6761 (The Origin of Shock Emission in Proto-Planetary Nebulae) The WF/PC-2 was used to make observations of the bi-polar proto-planetary nebula PK166-06D1 in order to study the origin of its shock emissions. The objective is to characterize the mass loss during the transition to planetary nebula status. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed WF/PC-2 6681 (Evolution of Pre-Main Sequence Circumstellar Nebulosity) The WF/PC-2 was used to make observations of V836-Tau to study the evolution of pre-main sequence circumstellar nebulosity. Initial HST results from imaging of disks around pre-main sequence stars provide compelling justification to observe a broader sample of young stars. The observations were completed as planned, and no anomalies were reported. 1.14 Completed WF/PC-2 6673 (Black Holes and Gas Disks in a Complete Sample of Radio Loud UGC Ellipticals) The WF/PC-2 was used to make observations of radio load, elliptical galaxy UGC-03695in order to search for possible black holes and gas disks. Hydrogen-Alpha and broad band observations were taken to determine the luminosity profiles and stellar populations (on the tens of parsec scale), and detect nuclear regions or disks of gas and dust. The observations were completed as planned, and no anomalies were reported. 1.15 Completed STIS/CCD 7476 (Sakurai's Nova-Like Object: Real-Time Monitoring of a Stellar Thermal Pulse) The Space Telescope Imaging Spectrograph (CCD) was used to make images of Sakurai's nova-like object in order to do real-time monitoring of a stellar thermal pulse. The eruption of Sakurai's nova-like variable in Sagittarius presents a "once-in-a-lifetime" opportunity for real-time observations of a star undergoing a final helium thermal pulse. A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.16 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.17 Completed NIC/1/2 7329 (The Nature of the Damped LyAlpha Absorbers -- A New Study of Young Galaxies) The Near Infrared Camera (Cameras One and Two) was used to make observations of the young galaxy Q1215+3322 in order to study the nature of its damped Lyman-Alpha Absorbers (DLAs). The DLA line systems, occasionally found in quasar spectra, provide one of only a few practical methods of detecting and studying significant numbers of galaxies at high redshifts. The observations were executed as scheduled, and no anomalies were noted. 1.18 Completed WF/PC-2 6458 (Imaging of QSO Host Galaxies at Z >2) The WF/PC-2 was used to make observations of the QSO/host galaxy FEIGE23. QSOs of redshift larger than 2 have been resolved with ~arcsecond resolution on 4m-class ground-based telescopes. Both line and continuum flux have been detected over a few arcseconds, but the nature of the extended emission, and its relation to the nuclear energy source is unclear. The observations were executed as scheduled, and no anomalies were noted. 1.19 Completed Two Sets of STIS/CCD 8080 (STIS Spectroscopic CCD Flatfielding -- Revised) The Space Telescope Imaging Spectrograph (CCD) was used to make CCD flat fielding calibration observations in the spectroscopic mode by make observations of object Feige23. The observations were completed as planned, and no anomalies were reported. 1.20 Completed WF/PC-2 7427 (Saturn's Rings) The WF/PC-2 was used to make imaging observations of both the east and west apsides of Saturn's rings. Saturn's rings are a unique dynamical laboratory and continue to present challenging problems which require additional data to attack. The observations were completed as planned, and no problems were reported. 1.21 Completed NIC/2/3 7453 (The SBF Hubble Diagram) The Near Infrared Camera (Camera Two and Three) was used to make deep NICMOS observations of galactic cluster Abell-0496 to determine its distance using the surface brightness fluctuation {SBF} method. HST is essential for its high spatial resolution which makes it possible to detect and remove globular clusters and background galaxies that contaminate the SBF measurement. The observations were executed as scheduled, and no anomalies were noted. 1.22 Completed WF/PC-2 6639 (Star Clusters and The Duration of Starbursts) The WF/PC-2 was used to make observations of the star cluster NGC-3690 in order to study the duration of its starbursts. Starbursts are a fascinating but poorly understood phenomenon that probably play a key role in galaxy evolution. The observations were completed as planned, and no problems were reported. 1.23 Completed STIS/CCD 7563 (Francis Cluster) The Space Telescope Imaging Spacetrograph (CCD) was used to make observations of the cluster Gal-Clus-2 (a.k.a., Francis Cluster) using various modes. A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.24 Completed NIC/2 7178 (Composition of the Saturn Bright Rings) The Near Infrared Camera (Camera Two) was used to make observations of the bright rings on the planet Saturn in order to study it structure. Groundbased spectroscopy established a quarter of a century ago that there must be substances other than water ice in Saturn's bright rings, as indicated by their UV reflectivity, which does not match that of pure ice. Speculation has ranged from silicate to carbonaceous material, but there remains no near-infrared evidence for anything but water ice. The observations were completed as planned, and no anomalies were noted. 1.25 Completed FGS 6881 (Determination of the Masses of the Hyades Binary VA351 and the Circumstellar Structure of the Supergiant VV Cep {HD 208816}) The Fine Guidance Sensors were used to make astrometric observations of the variable supergiant star VA351. A search with Hubble Space Telescope Fine Guidance Sensor 3 in TRANS mode of 36 probable Hyades cluster members in the magnitude range V=10--15 has yielded nine certain binary detections. The observations were executed as scheduled, and no anomalies were noted. 1.26 Completed WF/PC-2 6659 (A Search For Superplanets Around Weak T- Tauri Stars) The WF/PC-2 was used to make observations in the 1 micron band of the T-Tauri class star HBC426 in order to search for superplanets. Weak T-Tauri class stars are thought to have dissipated their circumstellar disks recently, possibly by forming planetary systems. The observations were completed as planned, and no anomalies were reported. 1.27 Completed WF/PC-2 6619 (High Resolution Images of QSO Lyman Alpha Absorbing Galaxies) The WF/PC-2 was used to make several images of galaxy/QSO-130112 in order to obtain high resolution observations of the Lyman Alpha absorbing galaxies within this field-of-view. The purpose is to directly identify the objects responsible for low and intermediate redshift QSO Lyman-Alpha systems. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 27 Successful : 27 Scheduled Reacquisitions : 18 Successful : 18 2.2 FHST Updates: Scheduled : 64 Successful : 63 As documented in HSTAR 6738, the roll delay updates at 298/23459Z failed due to FHST-1. Both subsequent full maneuver updates succeeded, as did the following acquisition. 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was reset ten times. Following the Gyro #6 failure described in the previous report, Gyro #6 was powered off at 296/1608Z. The SPC recovery load was patched to reflect a three gyro configuration at 296/1926Z. As directed by ROP NS-3, the NSSC-1 status buffer was dumped at 297/1923Z. A TTR was written for a 5-minute, 33-second unrecoverable loss of engineering data due to a failure to lock up in Mode 1. HSTAR 6739 describes a STIS MCE-1 reset which occurred at 299/031518Z while outside an SAA boundary and while the low voltage was on. ROP NS-5 was utilized to clear the engineering report. The STIS MAMA-1 will be recovered via SMS at the time of the next high voltage on commanding. 3.0 SIGNIFICANT FORTHCOMING EVENTS: The DF-224 vehicle clock is scheduled to have its most significant bit (MSB) set to a '1' this afternoon at approximately 204215Z. Since many ground software systems use I*4 (4 byte integer value) to represent the 32-bit vehicle clock, there was concern that the MSB would be treated as a sign bit. This could have resulted in clock times becoming negative when the 32nd bit is set. Over the past years several systems using the I*4 representation for vehicle time, have been modified. All systems, onboard and on the ground, were then checked to ensure that the rollover would cause no problems. In some cases, additional modifications were made. No problems with the rollover are anticipated. Continuation of normal science observations and calibrations. /DMH